Write Ln 3x + 4 Ln x - Ln 3xy as a single logarithm
Found 2 solutions by nerdybill, CharlesG2:
Answer by nerdybill(7384)
Answer by CharlesG2(834)

Write Ln 3x + 4 Ln x - Ln 3xy as a single logarithm
ln 3x + 4 ln x - ln 3xy
ln 3x + ln x^4 - ln 3xy
logarithmic rules: logb (mn) = logb (m) + logb (n)
logb (m/n) = logb (m) - logb (n)
in ln (natural log) case base b is e
ln (3x * x^4) - ln 3xy
ln 3x^5 - ln 3xy
ln (3x^5 / 3xy)
ln (x^4)/y